OpenStudy (anonymous):

(x^2 + 3x - 2)(x + 3)

OpenStudy (anonymous):

A. x3 + 6x2 + 7x - 6 B. x2 + 4x + 1 C. x3 + 3x2 + 1 D. x3 + 6x2 + 11x + 6

OpenStudy (anonymous):

@nightwill

OpenStudy (anonymous):

\[(x^2 + 3x - 2)(x + 3)\] Similarly, \[x(x^2+3x-2)+3(x^2+3x-2)=x^3+3x^2-2x+3x^2+9x-6\] \[=x^3+6x^2+7x-6\]

OpenStudy (anonymous):

Thank You very much.